If 10 9 electrons move out of a body to another body every second,how much time is required to get a total charge of 1 coulomb of the other body?

If 109 electrons move out of a body to another body every second,how much time is required to get a total charge of 1 coulomb of the other body?

Grade:12

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
According to the question, 109electrons move out in 1 sec.
So, the charge transferred per unit sec is equal to1.6\times 10^{-19} \times 10^9=1.6 \times 10^{-10} C
so, to transfer 1 C charge, the time taken would be equal to
\frac {1}{1.6 \times 10^{-10} C}=0.625\times 10^{10} sec
Regards